# REG4Rec: Reasoning-Enhanced Generative Model for Large-Scale Recommendation Systems

### v2

@@ 1. Introduction
Figure 1. Comparison of reasoning in GR: (a) GR with a limited reasoning process; (b) our REG4Rec enables diverse and reliable reasoning process to better capture the variety of users’ intents. :
@@ 2.1. Sequential Recommendation
2.2. Reasoning-Enhanced Generation
@@
Figure 2. REG4Rec introduces MPQ for generating flexible and reliable reasoning paths. During training, CRSS diversifies path combinations, while PARS selects consistent, confident paths. MSRA extends the reward horizon to capture long-term preferences and reduce noise caused by stochastic behavior. During inference, CORP prunes inconsistent paths, preventing the propagation of errors in the reasoning process. :
@@ 3.1. Problem Formulation
This section formalizes the GR task. Let $\mathcal{I}$ denote the set of all items. For a given user $u$, the historical interaction sequence is defined as $\mathcal{S}_{u}=(i_{1},i_{2},\ldots,i_{T})$, where $i_{t}\in\mathcal{I}$ is the item interacted with at time step $t$. The primary objective is to predict the user’s next interaction, $i_{T+1}\in\mathcal{I}$. Inspired by prior work (Rajput et al., 2023; Yang et al., 2025), we integrate the GR model into the retrieval stage of recommender systems. A common approach is to build an encoder-decoder GR model. The model takes $\mathcal{S}_{u}$ as input and estimates a distribution $p(i|\mathcal{S}_{u})$ over $i\in\mathcal{I}$. During inference, the model retrieves the top-$N$ items by sorting $p(i|\mathcal{S}_{u})$ in descending order.
@@ 3.2. Overview of REG4Rec
3.3. Flexible Reasoning Paths
@@ 3.3.1. MPQ: Moe-based Parallel Quantization Codebooks
Concretely, for an item $i$, its multi-modal features (e.g., textual, visual, and behavioral) are extracted with a pretrained model (Xing et al., 2025) and concatenated into a unified feature vector $v_{i}$. Let $M$ denote the number of parallel autoencoder experts $\{E_{r}\}^{M}_{r=1}$. Each expert $E_{r}$ is associated with a codebook $\mathcal{C}^{r}=\{z^{r}_{1},z^{r}_{2},\ldots,z^{r}_{D}\}$, where $z^{r}_{j}$ is a codeword and $D$ is the codebook size. Each expert operates independently, takes $v_{i}$ as input and projects it into a latent spaces $e^{r}_{i}=E_{r}(v_{i})$. The experts then select the most similar codeword in $\mathcal{C}_{r}$ to $e^{r}_{i}$, using its index to represent a semantic token for item $i$. Formally, the index $c^{i}_{r}$ which is selected by the $r$-th expert represent the semantic token, is calculated as:
@@ 3.4.2. PARS: Preference Alignment for Reasoning
Global Path Reward ($\gamma_{\text{path}}$): We evaluate path robustness with a progressive retrieval scheme that allows up to $d$ tokens to be incorrect, testing the model’s ability to maintain retrieval performance under partial errors. For a reasoning path of length $M$, we build a cumulative sum of codebook representation $z$ using only the MdM-d correct tokens, $z=\sum_{m=1}^{M-d}z^{m}$, and perform top-$N$ retrieval from the item pool to check whether the target item is still retrieved. The reward under $d$ errors is:
@@
In Eq. (14), the indicator function $\mathbb{I}(\cdot)$ return 1 if the target item is ranked within the top-$N$ retrieved candidates and 0 otherwise.
@@ 3.4.3. MSRA: Multi-Step Reward Augmentation Strategy
Additionally, We also extend the global path reward. Rather than requiring exact retrieval of the single target $i_{t+1}$, the augmented path reward is positive if the top-$N$ results include any of the $h$ candidate items $\{i_{t+1},\dots,i_{t+h}\}$. This relaxation encourages the model to accommodate broader user intents. Formally, let $W_{t}=\{i_{t+1},\dots,i_{t+h}\}$ represent the set of $h$ candidate items, and let $\text{Top-}N(z)$ be the top-$N$ retrieved items using the cumulative representation $z$ of the generated path:
@@
where $\mathbb{I}(\cdot)$ denotes an indicator function that returns 1 if any item in the candidate set is included in the top-$N$ results and 0 otherwise. By applying time-decayed weighting, MSRA prioritizes recent interactions while preserving sensitivity to long-term preferences, improving the robustness and diversity of the reasoning paths.
@@ 4.1. Experimental Setup
Implementation Details. This section summarizes our implementation settings. The training process is executed on a distributed TensorFlow (Abadi et al., 2016) platform, consisting of 10 parameter servers and 10 workers, each with 2 GPUs. To ensure a fair comparison with baselines, we configure the total number of transformer layers to 3, with one layer for the encoder and two layers for the decoder. The input dimension is 128 and the hidden size per transformer layer is 640. For MPQ, the hyperparameter $\alpha$ used in the training process of MPQ is set to 0.001. Because MPQ training is not the core focus and these weights have minimal impact once MPQ converges, we do not perform a sensitivity study on them. For flexible reasoning paths, MPQ uses 8 codebooks with 300 codewords each. The tokenizers process both text and image features for the industrial and public datasets. Text embeddings are generated using the pre-trained Qwen3-Embedding 7B model (Zhang et al., 2025a) and image embeddings are generated using the pre-trained Pailitao v8 model (Zhang et al., 2018), both with a dimensionality of 256. We set the clipping parameter $\epsilon$ to 0.15 to balance exploration and exploitation. In MSRA, the number of future steps is $h=3$ and the time-decay coefficient is $w=0.8$. In CORP, the self-reflection threshold is $\theta=0.06$. All other hyperparameters of these modules have been carefully tuned through sensitivity analysis, as detailed in Section 4.4.
@@ 4.4. Sensitivity Analysis (RQ3)
Figure 3. Sensitivity Analysis:
@@ 4.5. Reasoning Step Scaling (RQ4)
Figure 4. The performance of REG4Rec across different numbers of reasoning steps.:

### v3

@@ 1. Introduction
Figure 1. Comparison of reasoning in GR: (a) GR with a limited reasoning process; (b) our REG4Rec enables diverse and reliable reasoning process to better capture the variety of users’ intents.
$x$.
@@ 2.1. Sequential Recommendation
$2.$ Reasoning-Enhanced Generation
@@
Figure 2. REG4Rec introduces MPQ for generating flexible and reliable reasoning paths. During training, CRSS diversifies path combinations, while PARS selects consistent, confident paths. MSRA extends the reward horizon to capture long-term preferences and reduce noise caused by stochastic behavior. During inference, CORP prunes inconsistent paths, preventing the propagation of errors in the reasoning process.
$x$.
@@ 3.1. Problem Formulation
This section formalizes the GR task. Let $\mathcal{I}$ denote the set of all items. For a given user $u$, the historical interaction sequence is defined as $\mathcal{S}_{u}=(i_{1},i_{2},\ldots,i_{T})$, where $i_{t}\in\mathcal{I}$ is the item interacted with at time step $t$. The primary objective is to predict the user’s next interaction, $i_{T+1}\in\mathcal{I}$. Inspired by prior work (Rajput et al., 2023; Yang et al., 2025), we integrate the GR model into the retrieval stage of recommender systems. A common approach is to build an encoder-decoder GR model. The model takes $\mathcal{S}_{u}$ as input and estimates a distribution $p(i|\mathcal{S}_{u})$ over $i\in\mathcal{I}$. During inference, the model retrieves the top-NN items by sorting $p(i|\mathcal{S}_{u})$ in descending order.
@@ 3.2. Overview of REG4Rec
$3.$ Flexible Reasoning Paths
@@ 3.3.1. MPQ: Moe-based Parallel Quantization Codebooks
Concretely, for an item $i$, its multi-modal features (e.g., textual, visual, and behavioral) are extracted with a pretrained model (Xing et al., 2025) and concatenated into a unified feature vector $v_{i}$. Let $M$ denote the number of parallel autoencoder experts $\{E_{r}\}^{M}_{r=1}$. Each expert $E_{r}$ is associated with a codebook $\mathcal{C}^{r}=\{z^{r}_{1},z^{r}_{2},\ldots,z^{r}_{D}\}$, where $z^{r}_{j}$ is a codeword and $D$ is the codebook size. Each expert operates independently, takes $v_{i}$ as input and projects it into a latent spaces $e^{r}_{i}=E_{r}(v_{i})$. The experts then select the most similar codeword in $\mathcal{C}_{r}$ to $e^{r}_{i}$, using its index to represent a semantic token for item ii. Formally, the index $c^{i}_{r}$ which is selected by the $r$-th expert represent the semantic token, is calculated as:
@@ 3.4.2. PARS: Preference Alignment for Reasoning
Global Path Reward ($\gamma_{\text{path}}$): We evaluate path robustness with a progressive retrieval scheme that allows up to $d$ tokens to be incorrect, testing the model’s ability to maintain retrieval performance under partial errors. For a reasoning path of length $M$, we build a cumulative sum of codebook representation $z$ using only the MdM-d correct tokens, $z=\sum_{m=1}^{M-d}z^{m}$, and perform top-NN retrieval from the item pool to check whether the target item is still retrieved. The reward under $d$ errors is:
@@
In Eq. (14), the indicator function $\mathbb{I}(\cdot)$ return 1 if the target item is ranked within the top-NN retrieved candidates and 0 otherwise.
@@ 3.4.3. MSRA: Multi-Step Reward Augmentation Strategy
Additionally, We also extend the global path reward. Rather than requiring exact retrieval of the single target $i_{t+1}$, the augmented path reward is positive if the top-NN results include any of the $h$ candidate items $\{i_{t+1},\dots,i_{t+h}\}$. This relaxation encourages the model to accommodate broader user intents. Formally, let $W_{t}=\{i_{t+1},\dots,i_{t+h}\}$ represent the set of $h$ candidate items, and let $\text{Top-}N(z)$ be the top-NN retrieved items using the cumulative representation $z$ of the generated path:
@@
where $\mathbb{I}(\cdot)$ denotes an indicator function that returns 1 if any item in the candidate set is included in the top-NN results and 0 otherwise. By applying time-decayed weighting, MSRA prioritizes recent interactions while preserving sensitivity to long-term preferences, improving the robustness and diversity of the reasoning paths.
@@ 4.1. Experimental Setup
Implementation Details. This section summarizes our implementation settings. The training process is executed on a distributed TensorFlow (Abadi et al., 2016) platform, consisting of 10 parameter servers and 10 workers, each with 2 GPUs. To ensure a fair comparison with baselines, we configure the total number of transformer layers to 3, with one layer for the encoder and two layers for the decoder. The input dimension is 128 and the hidden size per transformer layer is 640. For MPQ, the hyperparameter $\alpha$ used in the training process of MPQ is set to 0.001. Because MPQ training is not the core focus and these weights have minimal impact once MPQ converges, we do not perform a sensitivity study on them. For flexible reasoning paths, MPQ uses 8 codebooks with 300 codewords each. The tokenizers process both text and image features for the industrial and public datasets. Text embeddings are generated using the pre-trained Qwen3-Embedding 7B model (Zhang et al., 2025a) and image embeddings are generated using the pre-trained Pailitao v8 model (Zhang et al., 2018), both with a dimensionality of 256. We set the clipping parameter $\epsilon$ to 0.15 to balance exploration and exploitation. In MSRA, the number of future steps is $h=3$ and the time-decay coefficient is $w=0.8$. In CORP, the self-reflection threshold is $\theta=0.06$. All other hyperparameters of these modules have been carefully tuned through sensitivity analysis, as detailed in Section $4.$
@@ 4.3. Ablation Study (RQ2)
Figure 3. Sensitivity Analysis
$x$.
@@ 4.5. Reasoning Step Scaling (RQ4)
Figure 4. The performance of REG4Rec across different numbers of reasoning steps.
$x$.